OpenStudy (anonymous):

how many tangents to the curve \[y=x^3-3x\]can be drawn from different points in the plane?

OpenStudy (anonymous):

\[\large \color{green}{y=x^3-3x}\]

OpenStudy (experimentx):

do the usual stuff .. take derivatives, find the critical points.

OpenStudy (anonymous):

the first derivative \[y'=0\\3x^2-3=0\\x=1,x=-1\]

OpenStudy (experimentx):

3x^2 - 3 = 0

OpenStudy (anonymous):

\[f(1)=-2,f(-1)=2\]

OpenStudy (anonymous):

but the didnt ask us to find critical points

OpenStudy (anonymous):

if we say there is a fixed \[x=x_0\] then the tangent line at this point is \[y-y_0=(3x_0^2-3)(x-x_0 )\\y=-x_0^3+3x+3xx_0^2-3x_0^3-3x+3x_0\\y=3x(x_0^2)+x_0(3-4x_0^2)\]

OpenStudy (anonymous):

@experimentX check the updated question

OpenStudy (anonymous):

the question was not complete...sry

OpenStudy (experimentx):

oh ... i thought you might be looking for tangent parallel to x-axis.

OpenStudy (anonymous):

but do u understand the new version of the problem

OpenStudy (experimentx):

no I don't ... there are infinite number of tangents ... isn't there some constraint?

OpenStudy (anonymous):

the concept of from different points is illusional

OpenStudy (experimentx):

?? what does that mean?

OpenStudy (anonymous):

i am trying to understand it...the least i am getting is that the tangents are drawn from this points and we cant use the same point twice

OpenStudy (experimentx):

the is only one tangent at one point ... the all tangents are unique.

OpenStudy (anonymous):

yes but now we are interste with other points on the graph ...all the points

OpenStudy (experimentx):

did you mean that tangent passing through any particular point (a,b) ??

OpenStudy (anonymous):

yes i mean thats how i wud say i understand it...the problem is not giving any such restrictionss....

OpenStudy (experimentx):